OpenStudy (anonymous):

V=43πr3 where r is the radius=10.5 cm.

OpenStudy (anonymous):

Where are you stuck?

OpenStudy (anonymous):

\[ V=\frac{4}{3}\pi r^3\\ r=10.5 \]Plug in \(r\) into the formula.

OpenStudy (anonymous):

on the fraction part 4/3 im bad with fractions